OpenStudy (anonymous):

find the value of x for which abcd is a parallelogram 25-3x=3x-17

OpenStudy (campbell_st):

add 3x to both sides 25 = 6x - 17 add 17 to both sides 42 = 6x hopefully you can solve it from here

OpenStudy (anonymous):

oh gosh thank you that is so obvious now

hero (hero):

Jordan.karst, you have to master the concept of like terms

hero (hero):

If you have 25 - 3x = 3x - 17 It should be obvious to recognize that 25 and 17 are like terms and 3x and 3x are like terms so putting like terms on the same side produces 25 + 17 = 3x + 3x Do you get that?

OpenStudy (anonymous):

yes thank you

hero (hero):

Like terms 1,2,3,4,5 are like terms because they all have the same no variables because they are integers 1x, 2x, 3x, 4x....are like terms because they all have the same variable, x 1y, 2y, 3y, 4y....27y are like terms because they all have the same variable, y 1xy, 2xy, 3xy, ....50xy are like terms because they all have the same variables xy 2x^2, 3x^2, 4x^2...are like terms because they all have the same variables x^2 and so on...